Find a particular solution of the given differential equation. Use a CAS as an aid in carrying out differentiations, simplifications, and algebra.

y^(4) + 2y'' + y = 1 cos x − 3x sin x
1
Expert's answer
2020-07-15T19:01:52-0400


Given D.E is

"y^{(4)} + 2y'' + y = \cos(x) − 3x\sin(x)"

Thus, auxilary equation will be

"m^4+2m^2+1=0\\
\implies (m^2+1)^2=0\\
\implies m=\pm i,\pm i"

Hence,

"C.F=(c_1+c_2x)\cos(x)+(c_3+c_4x)\sin(x)"

Now, from "C.F" we can guess the particular solution is

"y_p(x)=(ax+bx^2)\cos(x)+(cx+dx^2)\sin(x)"

Now, we have to find the value "a,b,c,d" .

Thus, find the the expression

"y_p^{(4)} + 2y_p'' + y_p"

And compare the coefficients of the above expression with "\cos(x) − 3x\sin(x)" we get

"y_p^{(4)} + 2y_p'' + y_p=\cos(x)(-24ax-8b+24c)+\sin(x)(-24a-24cx-8d)\\\implies 

-24a=0\\

-8b+24c=1\\

-24c=3\\

-24a-8d=0"

Thus,

"a=d=0,b=\frac{1}{4},c=\frac{1}{8}\\"


"\implies y_p(x)=\frac{x^2}{4}\cos(x)+\frac{x^3}{8}\sin(x)"